smith has 3 times as much money invested at 8% bonds as he has in stocks paying 9.5%. how much does he have invested in each if his yearly income from the investments is $6365 ?
Answer by checkley71(8403)

.08*3x+.095x=6,365
.24x=6,365
.335x=6,365
x=6,365/.335
x=19,000 amount invested @ 9.5%
proof
3*19,000*.08+.095*19,000=6,365
4,560+1,805=6,365
6,365=6,365